OpenStudy (anonymous):

can someone help me please? A town of 20,000 people in Germany generates 0.95 kg per person per day of MSW. A town of the same size in the US generates 1.9 kg per person per day of MSW. 1. How much MSW is generated in each town daily and weekly? 2. How many trucks would be needed to collect the waste twice weekly? The trucks each have a capacity of 4.5 metric tons. And operate 5 days per week. Assume that the trucks average two loads per day at 75% capacity. 3. If the towns recycle waste at 16% (Germany) and 23% (US), determine the mass and volume of MSW that enters the landfill daily

OpenStudy (anonymous):

1. Germany:1900kg daily 133000kg weekly. US: 3700kg daily 266000kg weekly.

OpenStudy (anonymous):

so far i have that part

OpenStudy (anonymous):

how do i solve for 2.?

OpenStudy (anonymous):

First determine how many kg(s) in 1 ton. and then multiply by 4.5 and then by .75

OpenStudy (anonymous):

1000kg in 1 ton.

OpenStudy (anonymous):

So 1000*4.5=4500*.75=3375

OpenStudy (anonymous):

That's how many kg in one load.

OpenStudy (anonymous):

ok gotcha. i'm sorry i'm not to good with this stuff

OpenStudy (anonymous):

It's fine. This is how my math class was in 4th grade. I hate advanced math.

OpenStudy (anonymous):

so that is what we get for problem 2 1375

OpenStudy (anonymous):

3375

OpenStudy (anonymous):

Not quite. See 3375 is for one load and 1 truck averages 2 loads a day.

OpenStudy (anonymous):

yah i never have been good with word problems

OpenStudy (anonymous):

okay so let me see...

OpenStudy (anonymous):

Hold on I have to change my baby real quick. hahaa. I'll be right back.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

if we have two loads we get 67500 as the final answer correct? let me know when you get back :)

OpenStudy (anonymous):

6750 is how much total one truck gets a day. So take Germany for the first part 133000kg is how much they produce weekly So divide by 2 since they will come twice weekly. 66500 then divide that by 6750 66500/6750=9.85 So you will need 10 trucks.

OpenStudy (anonymous):

alright got it

OpenStudy (anonymous):

Now that's just for Germany. For the US 266000kg is how much they produce weekly. So divide by 2 since they will come twice weekly. 133000 then divide that by 6750 133000/6750=19.7 So you will need 20 trucks.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

That's #2. Here's it completely written out. First determine how many kg(s) in 1 ton. and then multiply by 4.5 and then by .75 1000kg in 1 ton. So 1000*4.5=4500*.75=3375 That's how many kg in one load. 3375 is for one load and 1 truck averages 2 loads a day. 6750 is how much total one truck gets a day. So take Germany for the first part 133000kg is how much they produce weekly So divide by 2 since they will come twice weekly. 66500 then divide that by 6750 66500/6750=9.85 So you will need 10 trucks. For the US 266000kg is how much they produce weekly. So divide by 2 since they will come twice weekly. 133000 then divide that by 6750 133000/6750=19.7 So you will need 20 trucks.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

3. If the towns recycle waste at 16% (Germany) and 23% (US), determine the mass and volume of MSW that enters the landfill daily Germany:1900kg daily US: 3700kg daily 3. So for this one we do Germany: 1900*.16=304kg recycled material So 1900-304=1596kg which is the amount of waste. US: 3700*.23=851kg recycled material So 3700-851=2849kg which is amount of waste.

OpenStudy (anonymous):

see that's were i messed up i had 19000*.16=3040 lol

OpenStudy (anonymous):

Hahaa.

OpenStudy (anonymous):

got that part correct n ow

OpenStudy (anonymous):

Perfect.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem. I'll help anytime I can as much as I can.